The average of the ages of a man and his daughter is 34 years. If the respective ratio of their ages four years from now is 14:5, what is the daughter's present age?
(1) 12 (2) 18 (3) 10 (4) Cannot be determined (5) None of these

Answer by richwmiller(17219)

(m+d)/2=34
(m+4)/(d+4)=14/5
d=16 m=52
daughter is 16 now
(5) none of these
check
56/20=14/5
4*14/4*5=14/5
ok